The probability that a plant produces two or more seeds is 0.30. A sample of seven plants is observed. What is the probability that at most two plants produces two or more seeds?

0.5128

0.6471

0.4472

0.7373
Answer #1

Here, n = 7, p = 0.3, (1 - p) = 0.7 and x = 2
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X <= 2).
P(X <= 2) = (7C0 * 0.3^0 * 0.7^7) + (7C1 * 0.3^1 * 0.7^6) + (7C2 * 0.3^2 * 0.7^5)
P(X <= 2) = 0.0824 + 0.2471 + 0.3176
P(X <= 2) = 0.6471